OpenStudy (anonymous):

PLEASE HELP --> WILL GIVE MEDAL --> PLEASE EXPLAIN Task 3 Part 1: Create a scenario for an arithmetic sequence. For example: Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. _____________________________________ Part 2: Create a scenario for a geometric sequence. For example: Anthony goes to the gym for ______ minutes on Monday. Every day he _________his gym time by ____________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? __________________________________________________________ Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (benlindquist):

what are the word choices

OpenStudy (anonymous):

There is no wordbank

OpenStudy (anonymous):

but is my answer right?

OpenStudy (anonymous):

Jasmine practices the piano for _10_____ minutes on Monday. Every day she ____increases_______ her practice time by __15 minutes_______. If she continues this pattern, how many minutes will she practice on the 7th day? Use your scenario to write the function for the 7th term in your sequence using sequence notation. Monday- 10 minutes Tuesday - 25 minutes Wednesday - 40 minutes Thursday - 55 minutes Friday - 70 minutes Saturday - 85 minutes Sunday - 100 minutes She will practice piano for 100 minutes on the 7th day, Sunday. Function: An = a1 + d(n – 1) Day N = 10 +15(n-1) Day 7= 10+15 (7-1) Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10%________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. (I rounded the decimals here) Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes He will spend 29.282 minutes at the gym on the 5th day, Friday. Sn=a*1−r^n/1−r where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. S(5)=20 1-1.1^5 / 1-1.1 =122.102

OpenStudy (anonymous):

@I_luv_Sparkles

OpenStudy (anonymous):

im done bye

OpenStudy (anonymous):

ok, thx.......I guess............. :(

OpenStudy (benlindquist):

im checking

OpenStudy (anonymous):

kk

OpenStudy (benlindquist):

i think so. did this in 6th grade. medal please

OpenStudy (anonymous):

I'm sorry, but you need to explain to get a medal. How did you do this in 6th grade??? This is for freshmen.

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (dayakar):

your part 1 is correct it is in arithmetic progression

OpenStudy (dayakar):

part 2 you must take geometric progression monday let this is the first term a1= 20mi a2 = a1*r [here r = common ratio] suppose r=2 a2=20*2 =40mi a3= a2*2=40*2=80mi

OpenStudy (dayakar):

do u understand

OpenStudy (anonymous):

@dayakar is right @NiniNunu